OpenStudy (anonymous):

25a^2+5a-42 factor trinomial

OpenStudy (anonymous):

@Safari321

OpenStudy (safari321):

a(25a+5)-42 = (5a-6)(5a+7)

OpenStudy (anonymous):

\[25a^2+5a-42\] \[(5 a-6) (5 a+7)\] \[a = -\frac{ 7 }{ 5 }\] \[a = \frac{ 6 }{ 5 }\]

OpenStudy (anonymous):

so it is (5a-6)(5a+7)

OpenStudy (safari321):

yes some_someone is showing what the a could equal

OpenStudy (anonymous):

Yes :)